The arc length of a curve y = g(x) over interval as xsb is given as: sa 1+ (g(x))2 da Approximate the arc length of g(x)= eUse at least 4 decimal places in your calculations.
Answer #1

0 gSolution Given that, I JIH (gl (x)) da a 9tx) = e-X glu) =-e-n (glex)) ? Flee? -le- OsXzI N=5, h=b-a N 1- 5 th 0.22 .. fin) - Suomi du en 0 The values of tabled for x and X 0 0:4 0.8 1:2 1.6 2 Y 1.4142 1.3609 Yo y 11122 1:2358 42 1.0379 11w = 1.12422 Solution by composite simpsons Rule is 1.12422